Atoms of CO2 gas decay from upper to lower level with a spread in frequency is about 10kHz. Calculate the line shape if the broadening mechanism is found Doppler broadening ?
1
Expert's answer
2020-05-24T18:05:18-0400

For the Doppler broadening we can use the following equation to calculate the Gaussian lines hape:


"l_d=\frac{1}{\sigma_0}\sqrt{\frac{mc^2}{2\pi kT}}\text{ exp}\bigg[-\frac{-mc^2(\sigma-\sigma_0)^2}{2kT\sigma_0^2}\bigg]"

normalized to


"\int l_d\text{d}\sigma=1."